OpenStudy (bbb911):

This is Geometry: Simplify: square root of 36 divided by square root of 8.

OpenStudy (anonymous):

3/sqrt(2)

OpenStudy (angela210793):

=6/sqr8=6*sqr8/8=3sqr8/4

OpenStudy (amistre64):

\[\frac{\sqrt{36}}{\sqrt{8}}=\frac{6}{2\sqrt{2}}\] \[\frac{3}{\sqrt{2}}*\frac{\sqrt{2}}{\sqrt{2}}=\frac{3\sqrt{2}}{2}\]

OpenStudy (bbb911):

Thank you amistre :D

OpenStudy (amistre64):

yep :) and that was even without my eraser ;)

OpenStudy (bbb911):

i wish i was smart as you , im only smart in English , math my worst subject :o

OpenStudy (anonymous):

or 6*squaroot 8/8